# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

### v1

@@ 4.3. Latency Lower Bound
Given $\mathcal{O}$ the set of operation type and $R_{o}$ a count of available resources of type $o\in\mathcal{O}$, $OG$ the operation graph, $cp_{L}$ the critical path of $OG$, and $LO(o)$ the latency function for operation $o$, with $LO(o)\geq 1$. $N_{o}^{L}$ denotes the number of operations of type $o$ in $L$. We define $C_{L}=\sum_{n\in cp_{L}}LO(n)$ the critical path of $L$ weighted by latency of its operations. The minimal latency of a region $L$ is bounded by
$Lat_{R_{o}}^{L}\geq\max(C_{L},\max_{o\in\mathcal{O}}(\lceil N_{o}^{L}\times LO(o)/R_{o}\rceil))$
@@
1L1: for (i = 0; i < N; i++) S0: s[i] = 0;
2L2: for (i = 0; i < M; i++) { S1: q[i] = 0;
3 L3: for (j = 0; j < N; j++) {
4 S2: s[j]+=r[i]*A[i][j]; S3: q[i]+=A[i][j]*p[j];}}
Listing 1: Bicg
For example in Lst. 1, if we consider the sub-loop body composed of loops L3 (fully unrolled) and the statements $S2$ and $S3$ as straight-line code regions, we can calculate the critical paths for $S2$ and $S3$ as follows: For $S2$, the critical path is given by: $cp_{S2}=\max(L(+)+L(*),N\times(DSP_{+}+DSP_{*})/R_{o})$ with $DSP_{o}$ the number of DSP for the operation $o$. For $S3$, the critical path is determined by: $cp_{S3}=\max(L(+)\times\log(N)+L(*),N\times(DSP_{+}+DSP_{*})/R_{o})$, considering the possibility of a tree reduction. In this context, the critical path for the entire sub-loop body is the maximum of these two individual critical paths, expressed as $\max(cp_{S2},cp_{S3})$.
We now need to integrate loops and enable composition of latency bounds.
Loop Unrolling  Loop unrolling is an HLS optimization that aims to execute multiple iterations of a loop in parallel. Intuitively, for an unroll factor $UF\geq 1$, $UF$ replications of the loop body will be instantiated. If $TC_{l}\mod UF_{l}\neq 0$ then an epilogue code to execute the remaining $TC_{l}\mod UF_{l}$ iterations is needed.
@@ 4 S2: s[j]+=r[i]*A[i][j]; S3: q[i]+=A[i][j]*p[j];}}
Given a loop $l$ with trip count $TC_{l}$ and loop body $L$, unroll factor $UF\leq TC$ and $\mathcal{O}$ the set of operation type. Given available resources $R_{o}$ and latencies $L(o)\geq 1$ with $o\in\mathcal{O}$. Then the minimal latency of $l$ if executed in a non-pipelined fashion is bounded by:
$Lat_{R_{o}}^{l,S}\geq\lfloor TC/UF\rfloor*Lat_{R_{o}}^{L}$
The intuition of the proof of Theorem 4.2 is that there cannot exist a shorter critical path in the operation graph of the unrolled loop nest than in the non-unrolled loop body, that is there is no useless operation. This is therefore essential to require the absence of useless operations in the program, as otherwise a shorter critical path may be exposed after loop unrolling, making this theorem incorrect.
@@
Consider the scenario of loop $L0$ within Listing 1, which has been unrolled by a factor denoted as $UF\leq TC_{L0}$ where $TC_{L0}$ is the trip count of $L0$. The latency for one iteration of $S0$ is denoted as $Lat^{S0}_{R_{o}}>0$. In the absence of pipelining, the lower bound of the latency of this sub-loop body is: $\lfloor TC_{L0}/UF\rfloor\times Lat^{S0}_{R_{o}}$.
@@
Loop pipelining  Loop pipelining amounts to overlapping multiple iterations of the loop, so that the next iteration can start prior to the completion of the preceding one. The initiation interval (II) measures in cycles the delay between the start of two consecutive iterations. It is easy to prove our formula template accurately integrates the latency of pipelined loops with the $I$ operator. We compute the minimal II in function of the dependencies of the pipelined loop and the iteration latency of the operations of the statements during the NLP generation. Let $RecMII$ and $ResMII$ be the recurrence constraints and the resource constraints of the pipelined loop, respectively. We have $II\geq\max{(ResMII,RecMII)}$. $RecMII=\max_{i}\lceil\frac{delay(c_{i})}{distance(c_{i})}\rceil$ with $delay(c_{i})$ the total latency in dependency cycle $c_{i}$ and $distance(c_{i})$ the total distance in dependency cycle $c_{i}$. We suppose that $ResMII=1$, as we don’t know how the resource will be used by the compiler. Hence, if the loop is a reduction loop then the $II\geq\frac{IL_{reduction}}{1}$ with $IL_{reduction}$ the iteration latency of the operation of reduction. For a kernel like $for(j=0;j<N;j++)y[j]=y[j-2]+3;$ the $II\geq\lceil\frac{IL_{+}}{2}\rceil$.
Program latency lower bound
@@
We now focus on having the program latency lower bound, with resource constraints. This bound takes into account the limitations imposed by available resources, which can significantly affect the achievable performance. We assume here that the resources consumed are only consumed by the computing units and resource use by the computational unit of one operation can not be reused by the computational unit of another operation executing at the same time. We also assume that the compilers have implemented the pragma configuration given as input.
@@
For DSPs we suppose we have a perfect reuse, i.e., that the computation units for the same operation can be reused as soon as the computation unit is not in use. Under-estimating the resources used is fundamental to proving the latency lower bound, as otherwise another design that consumes less resources than predicted may be feasible, itself possibly leading to a better latency. The lower bound of a kernel is therefore the lower bound of the configuration such that the minimal resource consumption is less than or equal to the resource available.
Theorem 4.3 (Lower bound of the kernel under resource constraint).
@@
Given a loop body $L$, the set of statements $\mathcal{S}$, $\mathcal{O}$ the set of operation type, $N_{o}^{s}$ the number of operations $o\in\mathcal{O}$ for the statements $s$, $DSP_{o}$ the number of resources (DSPs) used for the operation $o\in\mathcal{O}$, $MCU_{o}^{s}$ the maximal computational unit the statement $s$ can use in parallel and the configuration of pragma which is constituted of the vectors $\vec{PV_{i}}$ for each loop. The configuration of pragma, which has the minimal lower bound and which respect $R_{used}^{min}=\sum_{o}\max_{s\in\mathcal{S}}(N_{o}^{s}\times DSP_{o}\times MCU_{o}^{s})\leq R_{o}$, is the lower bound of the kernel.
The intuition of the proof of the Theorem 4.3 is that $R_{used}^{min}$ is the minimal resource that the kernel can use for the configuration of pragma. This implies that if the $R_{used}^{min}>R_{o}$ the configuration over-utilizes resources. Hence, the lowest lower bound which does not over-utilize the resource is the lower bound of the kernel.

### v2

@@ 4.2.1. Latency Lower Bound
Given $R_{op}$ a count of available resources of type $op$, for each operation type, and $LO(op)$ the latency function for operation $op$, with $LO(op)\geq 1$. #$\#L(op)$ denotes the number of operations of type $op$ in $L$. We define $LO(\#OG_{cp}^{L})=\sum_{n\in cp}LO(n)$ the critical path weighted by latency of its operations. The minimal latency of a region $L$ is bounded by
$Lat_{R_{op}}^{L}\geq\max(LO(\#OG_{cp}^{L}),\max_{o\in op}(\lceil\#L(o)\times LO(o)/R_{o}\rceil))$
@@
1#pragma ACCEL unroll factor=uf
2L0: for (i = 0; i < N; i++)
3 S0: s[i] = 0;
4#pragma ACCEL pipeline
5L1: for (i = 0; i < M; i++) {
6 S1: q[i] = 0;
7#pragma ACCEL unroll
8 L2: for (j = 0; j < N; j++) {
9 S2: s[j]+=r[i]*A[i][j];
10 S3: q[i]+=A[i][j]*p[j];
11 }
12}
Listing 5: Bicg code: $s=r\times A;q=A\times p$
For example, in Lst. 5, if we consider the sub-loop body composed of loops L2 (fully unrolled) and the statements $S2$ and $S3$ as straight-line code regions, we can calculate the critical paths for $S2$ and $S3$ as follows: For $S2$, the critical path is given by: $cp_{S2}=\max(L(+)+L(*),N\times(DSP_{+}+DSP_{*})/R_{o})$ with $DSP_{o}$ the number of DSP for the operation $o$. For $S3$, the critical path is determined by: $cp_{S3}=\max(L(+)\times\log(N)+L(*),N\times(DSP_{+}+DSP_{*})/R_{o})$, considering the possibility of a tree reduction. In this context, the critical path for the entire sub-loop body is the maximum of these two individual critical paths, expressed as $\max(cp_{S2},cp_{S3})$.
We now need to integrate loops and enable the composition of latency bounds.
4.2.2. Loop Unrolling: partial unroll
Loop unrolling is an HLS optimization that aims to execute multiple iterations of a loop in parallel. Intuitively, for an unroll factor $UF\geq 1$, $UF$ replications of the loop body will be instantiated. If $TC_{l}\mod UF_{l}\neq 0$ then an epilogue code to execute the remaining $TC_{l}\mod UF_{l}$ iterations is needed.
@@ 4.2.2. Loop Unrolling: partial unroll
Theorem 4.5 (Minimal latency of a partially unrolled loop with factor UF).
Given a loop $l$ with trip count $TC_{l}$ and loop body $L$, and unroll factor $UF\leq TC$. Given available resources $R_{op}$ and latencies $L(op)\geq 1$. Given $L^{{}^{\prime}}$ the loop body obtained by replicating $UF$ times the original loop body $L$. Then the minimal latency of $l$ if executed in a non-pipelined fashion is bounded by:
$Lat_{R_{op}}^{l,S}\geq\lfloor TC/UF\rfloor\times Lat_{R_{op}}^{L^{{}^{\prime}}}$
1#pragma ACCEL parallel UF=4
2L1: for (i = 0; i < 8; i++) {
3 S0: a[i] = b[i] * c[i];
4}
Listing 6: Example before partial unrolling
1// explicitly unroll with UF=4
2L1: for (i = 0; i < 8; i+=4) {
3 S0: a[i] = b[i] * c[i];
4 S1: a[i+1] = b[i+1] * c[i+1];
5 S2: a[i+2] = b[i+2] * c[i+2];
6 S3: a[i+3] = b[i+3] * c[i+3];
7}
Listing 7: Example after partial unrolling
In Listing 6, the loop body denoted by $L$ is depicted in the rectangle. The loop L1 is subjected to a partial unrolling with a factor of 4, as specified by the directive #pragma ACCEL parallel UF=4 hence S0 have to be expanded and replicated four times. The explicit unrolling of the loop is illustrated in Listing 7, where S0, S1, S2 and S3 represent the loop body $L^{{}^{\prime}}$ of the Theorem B.10. The latency of the program in Listing 6 and 7 (which are equivalent) is greater or equal to $\lfloor\frac{TC}{UF}\rfloor$, in this case 2, multiply by the latency of the loop body $L^{{}^{\prime}}$. In this context, $L^{{}^{\prime}}$ comprises statements S0, S1, S2, and S3. These four statements are independent of each other and can be executed concurrently. Therefore, the lower bound on latency for $L^{{}^{\prime}}$ is equivalent to the latency of the multiplication operation.
Note this bound requires to build the operation graph for the whole loop body. This is straightforward for inner loops and/or fully unrolled loop nests, but impractical if the loop body contains other loops. We therefore define a weaker, but more practical, bound:
@@ 6 S3: a[i+3] = b[i+3] * c[i+3];
Given a loop $l$ with trip count $TC_{l}$ and loop body $L$, and unroll factor $UF\leq TC$. Given available resources $R_{op}$ and latencies $L(op)\geq 1$. Then the minimal latency of $l$ if executed in a non-pipelined fashion is bounded by:
$Lat_{R_{op}}^{l,S}\geq\lfloor TC/UF\rfloor*Lat_{R_{op}}^{L}$
@@
Consider the scenario of loop $L0$ within Listing 5, which has been unrolled by a factor denoted as $UF\leq TC_{L0}$ where $TC_{L0}$ is the trip count of $L0$. The latency for one iteration of $S0$ is denoted as $Lat^{S0}_{R_{o}}>0$. In the absence of pipelining, the lower bound of the latency of this sub-loop body is: $\lfloor TC_{L0}/UF\rfloor\times Lat^{S0}_{R_{o}}$.
Vitis allows to do a reduction with a tree reduction in logarithmic time with the option “unsafe-math”.
Theorem 4.7 (Minimal latency of a partially unrolled loop with factor UF for reduction loop with tree reduction).
Given a reduction loop $l$ with trip count $TC_{l}$ and loop body $L$, and unroll factor $UF\leq TC$. Given available resources $R_{op}$ and latencies $L(op)\geq 1$. Then the minimal latency of $l$, if executed in a non-pipelined fashion and the tree reduction is legal is bounded by:
$Lat_{R_{op}}^{l,S}\geq\lfloor TC/UF\rfloor\times Lat_{R_{op}}^{L}\times\lfloor\log_{2}(UF)\rfloor$
1L1: for (i = 0; i < 8; i++) {
2 c += a[i];
3}
Listing 8: Example of code demonstrating a reduction, where a tree reduction technique can be applied, as depicted in Figure 1.
Figure 1. Illustrating the concurrent execution depicted in Listing 12 through a tree reduction method, accomplished within logarithmic time.
Figure 1 represent the tree reduction of the Listing 12. Due to the reduction, we use a tree reduction in order to increase the parallelism of the reduction, and execute a loop of trip count $n=8$ in $\log_{2}(n)=3$ iterations.
4.2.3. Loop pipelining
@@
Loop pipelining amounts to overlapping multiple iterations of the loop, so that the next iteration can start prior to the completion of the preceding one. The initiation interval (II) measures in cycles the delay between the start of two consecutive iterations. It is easy to prove our formula template accurately integrates the latency of pipelined loops with the $I$ operator. We compute the minimal II in function of the dependencies of the pipelined loop and the iteration latency of the operations of the statements during the NLP generation. Let $RecMII$ and $ResMII$ be the recurrence constraints and the resource constraints of the pipelined loop, respectively. We have $II\geq\max{(ResMII,RecMII)}$. $RecMII=\max_{i}\lceil\frac{delay(c_{i})}{distance(c_{i})}\rceil$ with $delay(c_{i})$ the total latency in dependency cycle $c_{i}$ and $distance(c_{i})$ the total distance in dependency cycle $c_{i}$. We suppose that $ResMII=1$, as we do not know how the resource will be used by the compiler. Hence, if the loop is a reduction loop then the $II\geq\frac{IL_{reduction}}{1}$ with $IL_{reduction}$ the iteration latency of the operation of reduction. For a kernel like the Listing 9 the $II\geq\lceil\frac{IL_{+}}{2}\rceil$.
1for (j = 0; j < N; j++)
2 y[j] = y[j-2] + 3;
Listing 9: Demonstration of a code snippet showcasing a scenario where a loop pipelined with a dependency of distance 2 results in an initiation interval ($II$) that satisfies $II\geq\lceil\frac{IL_{+}}{2}\rceil$.
It follows a bound on the minimal latency of a pipelined loop:
Theorem 4.8 (Minimal latency of a pipelined loop with known II).
Given a loop $l$ with trip count $TC_{l}$ and loop body $L$. Given available resources $R_{op}$ and latencies $L(op)\geq 1$. Then the minimal latency of $l$ if executed in a pipelined fashion is bounded by:
$Lat_{R_{op}}^{l,P}\geq Lat_{R_{op}}^{L}+II*(TC_{l}-1)$
In Listing 5, loop $L1$ is pipelined. However, loop $L1$ serves is a reduction loop for statement $S2$, meaning that we must await the completion of writing $s[j]$ at iteration $i$ before we can read and write again at iteration $i+1$. Consequently, the initiation interval ($II$) is bounded by or greater than the latency ($IL_{+}$) of the addition operation, which constitutes the reduction operation.
4.2.4. Loop pipelining and unrolling
A loop $l$ with trip count $TC_{l}$ can be pipelined and partially unrolled with $UF<TC_{l}$, in this case there is loop splitting where the trip count of the innermost loop equal to the unroll factor and the trip count of the outermost loop equal to $\frac{TC_{l}}{UF}$.
Theorem 4.9 (Minimal latency of a pipelined loop with known II and partially unrolled).
Given a loop $l$ with trip count $TC_{l}$, partially unrolled by an unroll factor $UF<TC_{l}$ and a loop body $L$. Given available resources $R_{op}$ and latencies $L(op)\geq 1$. Given $L^{{}^{\prime}}$ the loop body obtained by replicating UF times the original loop body $L$. Then the minimal latency of $l$ if executed in a pipelined fashion is bounded by:
$Lat_{R_{op}}^{l,P}\geq Lat_{R_{op}}^{L^{{}^{\prime}}}+II*(\frac{TC_{l}}{UF}-1)$
4.2.5. Non-Parallel, Non-Pipelined Loops
We continue with a trivial case: if the loop is not optimized by any directive (including any automatically inserted by the compilers), i.e., not parallelized nor pipelined, then every next iteration of the loop starts only after the end of the prior iteration.
Definition 4.10 (Lower bound on latency of a non-parallel, non-pipelined loop under resources constraints).
Given a loop $l$ with trip count $TC_{l}$ which is neither pipelined nor parallelized, that is, iteration $i+1$ starts after the full completion of iteration $i$, for all iterations. Given $Lat_{R_{op}}^{L}$ the minimal latency of its loop body. Then
$Lat_{R_{op}}^{l}\geq TC_{l}*Lat_{R_{op}}^{L}$
4.2.6. Coarse-Grained parallelization
Coarse-grained parallelization is a performance enhancement technique involving the unrolling of a loop which iterates a loop body not fully unrolled i.e., containing at least a pipelined loop or a loop executed sequentially. It is therefore impossible to do a coarse-grained parallelization with a reduction loop because the $n$ sub loop body are dependent on each other.
It follows a bound on the minimal latency of a coarse-grained unrolled loop:
Theorem 4.11 (Minimal latency of coarse-grained unrolled loop).
Given a loop $l$, which is not a reduction loop, with trip count $TC_{l}$, an unroll factor $UF\leq TC_{l}$ and $L$ the loop body iterated by the loop $l$ with a latency lower bound $Lat^{L}_{R_{op}}$. Given available resources $R_{op}$ and latencies $L(op)\geq 1$. Given $L^{{}^{\prime}}$ the loop body obtained by replicating $UF$ times the original loop body $L$. Then the minimal latency of $l$ if executed in a non-pipelined fashion is bounded by:
$Lat_{R_{op}}^{l,S}\geq\lfloor TC/UF\rfloor\times Lat_{R_{op}}^{L^{{}^{\prime}}}$
4.2.7. Program latency lower bound under resource constraints
@@
We now focus on the latency lower bound of a program, under resource constraints. This bound takes into account the limitations imposed by available resources, which can significantly affect the achievable performance. We assume here that the resources consumed are only consumed by the computing units and resource use by the computational unit of one operation can not be reused by the computational unit of another operation executing at the same time. We also assume that the compilers have implemented the pragma configuration given as input.
@@
For DSPs we suppose we have a perfect reuse i.e., that the computation units for the same operation can be reused as soon as the computation unit is not in use. Under-estimating the resources used is fundamental to proving the latency lower bound, as otherwise another design that consumes less resources than predicted may be feasible, itself possibly leading to a better latency.
Theorem 4.12.
@@
Given a loop body $L$, the set of set of statements $\mathcal{S}_{seq}$ non executed in parallel, #$\#L_{op}^{s}$ the number of operations $op$ for the statements $s$, $DSP_{op}$ the number of resources (DSPs) used for the operation $op$, $MCU_{op}^{s}$ the maximal number of computational units the statement $s$ can use in parallel at any given time, and the configuration of pragma $\vec{PV_{i}}$ for each loop. The minimal number of resource (DSPs) consumed, $R_{used}^{min}$, by $L$ for the pragma configuration is the sum, for each operation, of the maximum number of DSPs used in parallel by a statement. This corresponds to:
$R^{min}_{used}=\sum_{op}\max_{\mathcal{S}\in\mathcal{S}_{seq}}(\sum_{s\in\mathcal{S}}\#L_{op}^{s}\times DSP_{op}\times MCU_{op}^{s})$
Given a program and the available resource of DSP $DSP_{avail}$, if $R^{min}_{used}<DSP_{avail}$ the lower bound is valid and the program does not over-utilize the resources.
In Listing 5, statements S0 and S1 entail no DSP usage as they solely serve for initialization purposes. S2 and S3 undergo unrolling N times and operate independently, enabling parallel execution. Particularly, S3 necessitates a tree reduction since it acts as a reduction for loop $L2$.
Assuming 1 DSP for addition and 2 for multiplication (denoted as $DSP_{+}$ and $DSP_{*}$ respectively), the maximum number of computational units utilized by S2 and S3 for both addition and multiplication is N. Therefore, $MCU_{+}^{S2}=N$, $MCU_{*}^{S2}=N$, $MCU_{+}^{S3}=N$, and $MCU_{*}^{S3}=N$. Furthermore, each statement entails only one addition and one multiplication, hence #$\#L_{+}^{S2}=1$, #$\#L_{*}^{S2}=1$, #$\#L_{+}^{S3}=1$, and #$\#L_{*}^{S3}=1$.
Consequently, the minimum number of DSPs utilized by Listing 5 is $N\times(1+2)\times 2$.
4.2.8. Memory transfer
AMD/Xilinx Merlin manages automatically the memory transfer. The memory transfer and computation are not overlap (no dataflow) hence the latency is the sum of the latency of computation and communication. We assume that for each array the contents of the array are in the same DRAM bank.
Theorem 4.13 (Lower bound of the memory transfer latency for an array).
Given a loop body $L$, the set of array $\mathcal{A}$, an array $a\in\mathcal{A}$, and $LAT^{mem}_{a}$ the latency to transfer the array $a$ from off-chip to on-chip (inputs) and from on-chip to off-chip (outputs). $\forall a\in\mathcal{A},LAT^{mem}_{a}\geq(\mathds{1}_{a\in V_{O}^{L}}+\mathds{1}_{a\in V_{I}^{L}})\times footprint_{a}/max\_burst\_size$. With $\mathds{1}_{cond}=1$ if $cond$ = true else 0.
Within Listing 5, the matrix A is solely read and possesses a footprint of $N\times M\times 32$ bits. Given the FPGA’s limitation to process a maximum of 512 bits per cycle, the latency required for transferring A amounts to $\frac{N\times M}{16}$ cycles.
Theorem 4.14 (Lower bound of the memory transfer latency).
Given a loop body $L$, the set of array $\mathcal{A}$, the number of cycles to transfer the array $a$ is bounded by $\max_{a\in\mathcal{A}}(\mathds{1}_{a\in V_{O}^{L}}+\mathds{1}_{a\in V_{I}^{L}})\times footprint_{a}/max\_burst\_size$.
In Listing 5, when transferring all arrays before any computation, we can concurrently transfer each array (assuming they are in different DRAM banks). The primary bottleneck arises from the transfer of array $A$, which incurs a latency of $\frac{N\times M}{16}$ cycles. Upon program completion, arrays $s$ and $q$ need to be transferred back to DRAM, with a minimal latency of $\max(\frac{N}{16},\frac{M}{16})$ cycles each. Thus, the total communication latency amounts to $\frac{N\times M}{16}+\max(\frac{N}{16},\frac{M}{16})$ cycles.
4.3. Summary
By composing all the theorems, this allows us to end up with the final latency lower bound of the program which is presented in theorems B.20 for the computation and B.21 for the computation and communication.
Theorem 4.15 (Computation Latency Lower bound of a Program).
Given available resource $DSP_{avail}$, the properties vector $\vec{PV}_{i}$ for each loop and a program which contains a loop body $L$. The properties vector allows to give all the information concerning the trip counts and the II of the pipelined loops and to decompose the loop body $L$ with a set of loops $\mathcal{L}^{non\;reduction}_{L}$ potentially coarse-grained unrolled with $\forall l\in\mathcal{L}_{L},UF_{l}$ and a set of reduction loops executed sequentially $\mathcal{L}^{reduction}_{L}$ which iterates a loop body $L_{pip}$. By recursion the loop body $L_{pip}$ contains a pipelined loop $l_{pip}$ which iterate a loop body $L_{fg}$ fully unrolled. The loop body $L_{fg}$ contains operations which can be done in parallel with a latency $Lat^{L_{par}}_{R_{op}}$ and operations which are reduction originally iterated by the loops $\mathcal{L}^{reduction}_{L_{fg}}$ with a latency $Lat_{L_{seq}}$.
The computation latency lower bound of $L$, which respected $DSP^{min}_{ued}\leq DSP_{avail}$, executed with tree reduction is:
$Lat^{L}_{R_{op}}\geq\prod_{l\in\mathcal{L}^{par}_{L}}\frac{TC_{l}}{UF_{l}}\times\prod_{l\in\mathcal{L}^{reduction}_{L}}TC_{l}\times Lat^{L_{pip}}_{R_{op}}$
with
$Lat^{L_{pip}}_{R_{op}}=(Lat^{L_{fg}}_{R_{op}}+II\times(\frac{TC_{l_{pip}}}{l_{pip}\_UF}-1))$ and $Lat^{L_{fg}}_{R_{op}}=Lat_{L_{par}}+Lat_{L_{seq}}\times\prod_{l\in\mathcal{L}^{reduction}_{L_{fg}}}\frac{TC_{l}}{UF_{l}}\times\log_{2}(UF_{l})$.
In Listing 5, let’s denote the loop body with loop $L0$ and statement $S0$ as $LB0$, and the one with loops $L1$ and $L2$, along with statements $S1$, $S2$, and $S3$, as $LB1$.
For $LB0$, where the loop is unrolled with a factor $uf$ and no dependency exists between loop iterations, the latency is bounded by $\frac{N}{uf}\times 1$, assuming the initialization takes at least one cycle.
Regarding $LB1$, loop $L2$ is fully unrolled, with no dependencies between iterations for statement $S2$. As there is no dependency between $S2$ and $S3$, they can be executed in parallel, and the latency of the unrolled part equals the critical path, which is the maximum between the unrolled parts of $S2$ and $S3$ in this scenario. However, for statement $S3$, despite the absence of dependencies among multiplication operations enabling parallel execution, a tree reduction is involved. Given the pipelined nature of loop $L1$, functioning as a reduction for statement $S2$, we observe $II>IL_{+}$, where $IL_{+}$ denotes the latency of the reduction operation. Consequently, the latency of $LB1$ is bounded by $(\max(IL_{+}+IL_{*},IL_{+}\log_{2}(N)\times IL_{+})+IL_{+}\times(M-1))$.
Theorem 4.16 (Latency lower bound of a program optimized with Merlin pragmas).
Given available resource $DSP_{avail}$ and a program which contains a loop body $L$ with a computation latency $Lat_{L}^{computation}$ and a communication latency $Lat_{L}^{communication}$.
The lower bound for $L$ which respected $DSP^{min}_{ued}\leq DSP_{avail}$ and where the computation and communication can not be overlap is:
$Lat_{L}=Lat_{L}^{computation}+Lat_{L}^{communication}$